OpenStudy (anonymous):

Mental Arithmetic Are there any tricks to solve the following problem, mentally. Not Using a calculator 3e-9 / 0.0015

OpenStudy (rational):

\[\large \dfrac{3e^{-9}}{0.0015}\] like that ?

OpenStudy (rational):

or perhaps this \[\large \dfrac{3\times 10^{-9}}{0.0015}\]

OpenStudy (anonymous):

yeah thats right, how can u do that mentally?

OpenStudy (anonymous):

3 x 10^-9 / 0.0015

OpenStudy (anonymous):

@rational

OpenStudy (vivek3461):

The numerator is given in scientific notation. You convert the denominator in scientific notation. It will be easier to do without a calculator.

OpenStudy (anonymous):

0.0015 is 15*10^(-4) You have 3*10^(-9) divided to 15 * 10^(-4). Divide 3 to 15 and 10*(-9) to 10^(-4) individually and multiply the two results. Thus you have 1/5 which is 0.2 - multiplied with 10^(-5). 10^(-5) means just adding zero 5 times after the point and before the number, so the answer is : 0.000002 or 2*10^(-6) I love mental math.

OpenStudy (phi):

I would say 1.5 divides into 3 to give us 2 to get 1.5 in the bottom we multiply by 10^3. to compensate, we also multiply by 10^-3 \[ \dfrac{3\times 10^{-9}}{0.0015} = \\ \frac{3 \cdot 10^{-9}}{1.5 \cdot 10^{-3} }\] thus we expect the answer to be 2 * 10^some_exponent -9 - -3 = -9 +3 = -6 so \[ 2 \cdot 10^{-6} \] which you can do in your head if you are good at this sort of thing.